OpenStudy (anonymous):

Zoran has two toy wheels. The smaller wheel having a radius of 12 cm revolves four times to cover a certain distance. The larger wheel revolves three times to cover the same distance. What is the radius of the larger wheel? A.24 cm B.16 cm C.18 cm D.48 cm

OpenStudy (anonymous):

@Japs

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

is it 12*3

OpenStudy (anonymous):

Try setting up equations with the circumference...

OpenStudy (anonymous):

Circumference = 2 * pi * radius So the circumference of the first wheel is 2 * pi * 12 = 24 * pi

OpenStudy (anonymous):

whats pi

OpenStudy (anonymous):

\[\pi \]

OpenStudy (anonymous):

It's part of the formula for circumference, area of a circle, etc. It's about 3.14, but it's an irrational number, so the decimals go on forever... but you don't need to use the decimal form here... just leave it as " π" Circumference = 2 * π * radius

OpenStudy (anonymous):

is it 75.36

OpenStudy (anonymous):

JakeV8 plz help

OpenStudy (anonymous):

Have you taken algebra, or is this before algebra?

OpenStudy (anonymous):

oops, this is precalc, I see from the topic :)

OpenStudy (anonymous):

75.36 is what you get if you multiply 2 * 3.14 * 12, but don't multiply it with 3.13, just treat "pi" like a variable... so you get circumference as 24 * π

OpenStudy (anonymous):

so the answer is 24

OpenStudy (anonymous):

So that's how far around the wheel is, and it's how far the wheel travels in one rotation. So four rotations would be 4 * (24 π) = 96 π And you know that 4 rotations of the first wheel is equal to 3 rotations of the other wheel. So 96 π = 2 * π * (radius of larger wheel) or 96 π = 2 * π * radius (just to make it easier to write) Now you simplify that expression to solve for the radius... The " π " cancels out...

OpenStudy (anonymous):

so its 24 right

OpenStudy (anonymous):

oops, I made a mistake... forgot to multiply by 3 for the 2nd wheel taking 3 rotations...

OpenStudy (anonymous):

the answer is 16 right

OpenStudy (anonymous):

96 π = 3 * (2 π * radius) = 6π * radius) (96 π ) / (6 π) = radius 16 = radius

OpenStudy (anonymous):

yes, you got it. sorry 'bout that error..

OpenStudy (anonymous):

John deposited $20,000 for 2 years with a friend who gave him 12% simple interest. How much more interest will he receive if he deposits the same sum in a bank that offers 12% compound interest? $288 $290 $240 $200

OpenStudy (anonymous):

how about that one

OpenStudy (anonymous):

whoa, maybe you can repost this in a new question? That way others can see it... I would have to go look up compound interest formulas...it's been awhile :)

OpenStudy (anonymous):

okay